# Vibe coding: programming through conversation with artificial intelligence

### v1

@@ 4.5 Limitations
We note the novelty of vibe coding, which can be said to have begun in earnest in February 2025 with the publication of the Karpathy canon. This meant that our corpus of videos was limited to an initial ~35 hours of vibe coding content, ~8.5 hours of which represented high quality videos that we analysed. This is unquestionably quite a small corpus, and future work is necessary to expand our analysis. Moreover, as outlined from the outset in Section 1, vibe coding is a rapidly evolving, emergent, and negotiated practice, which means that our data bears witness to vibe coders merely beginning to undertake such exploration and negotiation. There is much still to learn about how the practice changes as vibe coders gain experience and develop new rhythms in the vibe coding workflow and as IDEs evolve to support them.
@@
Finally, none of the selected videos depicted non-expert end-users (in particular, non-programmers). So, while our findings show how experienced coders can apply their expertise to address many of the challenges of vibe coding workflows, we do not know how non-programmers might meet these challenges and how they create barriers to effective vibing. Critically, we have studied the role of expertise in vibe coding in some detail (e.g., Section 3.7), and concluded that programming and AI expertise is not only important, but essential for successful vibe coding. But we can only claim that this is true insofar as the vibe coder is in possession of said expertise, because we have not studied non-experts. Thus, research is needed to understand what kind of vibe coding – perhaps an altogether different activity, despite bearing superficial resemblances – might be practiced by non-experts.
@@ 5 Conclusion
This study is the first empirical analysis of vibe coding, an emergent programming paradigm where developers primarily author and edit code by interacting with code-generating AI through natural language prompts rather than direct code manipulation. Through framework analysis of curated think-aloud videos from YouTube and Twitch, we examined how programmers form goals, conduct workflows, and deploy expertise when engaging in this novel form of programming.

### v2

@@ 4.5 Limitations
We note the novelty of vibe coding, which can be said to have begun in earnest in February 2025 with the publication of the Karpathy canon. This meant that our corpus of videos was limited to an initial ~35 hours of vibe coding content, ~8.5 hours of which represented high quality videos that we analysed. This is unquestionably quite a small corpus, and future work is necessary to expand our analysis. Moreover, as outlined from the outset in Section 1, vibe coding is a rapidly evolving, emerging, and negotiated practice, which means that our data bears witness to vibe coders merely beginning to undertake such exploration and negotiation. There is much still to learn about how the practice changes as vibe coders gain experience and develop new rhythms in the vibe coding workflow and as IDEs evolve to support them.
@@
Finally, none of the selected videos depicted non-expert end-users (in particular, non-programmers). So, while our findings show how experienced coders can apply their expertise to address many of the challenges of vibe coding workflows, we do not know how non-programmers might meet these challenges and how they create barriers to effective vibe coding. Critically, we have studied the role of expertise in vibe coding in some detail (e.g., Section 3.7), and concluded that programming expertise is not only important, but essential for successful vibe coding. But we can only claim that this is true insofar as the vibe coder is in possession of said expertise, because we have not studied non-experts. Thus, research is needed to understand what kind of vibe coding – perhaps an altogether different activity, despite bearing superficial resemblances – might be practiced by non-experts.
@@ 5 Conclusion
This study is the first empirical analysis of vibe coding, an emerging programming paradigm where developers primarily author and edit code by interacting with code-generating AI through natural language prompts rather than direct code manipulation. Through framework analysis of curated think-aloud videos from YouTube and Twitch, we examined how programmers form goals, conduct workflows, and deploy expertise when engaging in this novel form of programming.
@@ Appendix A Video sources
Appendix B Analysis framework
Top Level Category Subcategories Elaboration
Goals a. What do they build? i. What applications did vibe coders build?
Intentions a. Initial intention i. How is the intention for an app formed and refined in vibe coding?
b. Is intention refined? How and why? i. What AI workflow resources (e.g., exploration, seeking alternatives) support and refine this intent?
Workflow a. Stages of workflow i. What are the stages of the vibe coding workflow?
b. Portfolio of tools and technologies i. How do different tools work together in the context of vibe coding?
b. Explicit prompting strategies i. What are the prompting strategies in vibe coding?
c. Granularity of prompts – high/low level i. What is the granularity of prompts and iterations?
d. Single/multi-objective prompts i. Do users address multiple issues at once or handle one problem at a time?
e. What prompting modes are used – voice/text/images? i. What is the role of different modes (voice, text, screenshot, etc.) in vibe coding?
Debugging a. What debugging strategies are applied? i. What debugging strategies are employed in vibe coding?
Challenges a. What challenges (besides code bugs) are encountered in achieving goals? i. What challenges and barriers are encountered during vibe coding?
b. How are they addressed? i. What strategies are adopted to address these challenges?
ii. In what ways do experts deploy their expertise?
b. What kinds of expertise are used? i. Is the relevant expertise strictly developer knowledge, or does it include understanding system requirements, end-to-end design, and the broader technical ecosystem?
c. When do they switch to manual work? i. When do practitioners transition from vibe coding to manual editing (code, text, image, artifact)?
Definition of vibe coding and performance a. How do programmers define vibe coding? i. When is an activity considered vibe coding versus simply using AI tools?
b. How does performance for YouTube/Twitch skew production? i. How does the performative aspect of vibe coding (e.g., YouTube demonstrations) influence the practice and perception of generative AI in programming?